# Neural Radiance Fields in Medical Imaging: Challenges and Next Steps

### v3

@@ Abstract
Neural Radiance Fields (NeRF), as a pioneering technique in computer vision, offer great potential to revolutionize medical imaging by synthesizing three-dimensional representations from the projected two-dimensional image data. However, they face unique challenges when applied to medical applications. This paper presents a comprehensive examination of applications of NeRFs in medical imaging, highlighting four imminent challenges, including fundamental imaging principles, inner structure requirement, object boundary definition, and color density significance. We discuss current methods on different organs and discuss related limitations. We also review several datasets and evaluation metrics and propose several promising directions for future research.
@@ 1 Introduction
Neural Radiance Fields (NeRFs) significantly advance traditional 3D reconstruction methods in several key aspects [1]. Unlike conventional techniques that often rely on geometric or voxel-based approaches, NeRFs utilize deep learning to create highly detailed and photorealistic 3D models from a set of 2D images [2]. Compared with traditional reconstruction methods, this approach allows for rendering complex scenes with intricate light and shadow interplay, which traditional methods struggle to capture accurately. NeRFs excel in handling occlusions and view-dependent effects, producing more continuous and smoother representations of 3D spaces. Additionally, NeRFs require fewer manual interventions and assumptions about scene geometry, leading to more automated and scalable 3D reconstructions [3]. This capability is particularly advantageous in medical imaging, where precise and detailed visualization of anatomical structures is crucial, offering enhanced diagnostic and surgical planning tools over traditional methods [4].
@@
By employing deep learning techniques to synthesize highly detailed and photorealistic 3D models from sets of 2D medical images, such as CT scans or MRIs, NeRFs provide an unprecedented level of detail and accuracy in visualizing complex anatomical structures [1]. This technique surpasses traditional 3D reconstruction methods by effectively handling intricate light interactions and view-dependent phenomena, thus providing clearer, more continuous representations of internal body structures [7]. Furthermore, NeRFs can potentially minimize the need for multiple or high-resolution scans that expose patients to excessive ionizing radiation [5]. Although the NeRFs-based method has yielded surprising results in traditional 3D image reconstruction, it cannot be directly applied to medical image reconstruction due to the unique internal details of medical images [8, 3].
This paper aims to highlight the significant motivation and challenges associated with implementing NeRFs in medical imaging. We have systematically categorized existing methods based on organ types, acknowledging that the distinct characteristics of various organ structures, such as sparsity, overlap, and visibility, pose considerable hurdles to achieving comprehensive morphological reconstructions. These obstacles often lead to incomplete 3D models, potentially omitting crucial details (see Figure 1 for an example). Then, we explore the future research opportunities that could contribute to overcoming these obstacles and advancing the field.
The key contributions of this work are outlined as follows:
  • We outline the key motivation and challenges associated with applying NeRFs to medical imaging, identifying the main hurdles that must be overcome.
@@
  • Furthermore, we explore potential future research avenues, emphasizing how using NeRFs in medical imaging can revolutionize diagnostic processes, surgical planning, and medical education. This represents a significant advancement in medical visualization and patient care, promising to enhance accuracy and efficacy of treatments.
The remaining content of this paper is organized as follows. After introducing the motivation and challenges of NeRFs in medical imaging in Section 2, we describe the medical image NeRF applications for different organs and discuss the limitations of the current methods in Section 3. In Sections 4, we summarize the public datasets for NeRFs in medical images. We discuss the future directions in Section 5 and conclude the paper in Section 6.
@@ 2.1 Motivation
(1) Minimizing Radiation Exposure. NeRFs in medical imaging can potentially reduce the necessity for multiple or high-resolution scans that subject patients to high levels of ionizing radiation. By generating 3D representations from a sparse collection of 2D images, these methods can decrease the frequency of exposures, lowering radiation doses and diminishing the risks associated with radiation exposure. This technique not only reduces the risk of radiation-related health conditions but also alleviates issues related to prolonged patient immobility during imaging sessions.
(2) Reducing Time Consumption. Compiling multiple scans into a coherent 3D representation is time-intensive, posing significant challenges in time-sensitive medical situations. NeRFs in medical imaging streamline this process by enabling the quick creation of 3D models from a limited number of images, thereby accelerating data collection and image reconstruction. This rapid processing is particularly advantageous in emergency medicine settings, where efficiency can directly impact patient care and outcomes.
(3) Lowering Imaging Costs. Achieving detailed 3D images with conventional CT scans requires many X-ray projections, each with a fine slice thickness, which can be costly. NeRFs in medical imaging address this issue by efficiently producing detailed 3D models from smaller images, potentially reducing the financial burden. The capability of NeRFs in medical imaging to deliver precise volumetric reconstructions with fewer data inputs suggests a decrease in the reliance on expensive imaging resources, offering a cost-effective alternative to traditional imaging methods.
(4) Reducing Motion-related Artifacts. Medical imaging reconstructions typically involve multiple scan acquisitions, causing lengthy sessions that necessitate patient stillness, thereby increasing the risk of motion artifacts. Specifically, Magnetic Resonance Imaging (MRI) sequences are highly susceptible to motion-related artifacts caused by breathing and cardiac pulsation, potentially leading to inaccuracies in 3D reconstruction. Such inaccuracies are critical in clinical settings, where precise anatomical measurements are essential for accurate treatment planning and patient monitoring.
@@ 2.2 Challenges
(1) Different Imaging Principles. Fundamental imaging principles in consumer electronics and medical imaging are different. For example, X-ray while color-based (RGB) NeRF methods have seen considerable development. Yet, their direct application to X-ray imaging may yield suboptimal outcomes due to inherent differences between visible light and X-ray modalities. Unlike visible light imaging, which captures external features by reflecting off an object’s surface, X-ray imaging penetrates the object and attenuates, creating images that primarily reveal internal structures. These internal details are crucial for 3D X-ray reconstruction, highlighting the need for tailored NeRF algorithms to effectively accommodate unique characteristics of X-ray imaging [9].
(2) Need for Detailed Inner Structures. Medical images often require high levels of detail to represent intricate inner structures of the body accurately. NeRFs, which synthesize 3D models from 2D images, must be capable of capturing these details to be clinically useful. However, accurately rendering the complex internal anatomy, such as the fine distinctions between different types of tissues or the subtle anomalies indicative of early-stage diseases, can be challenging. This is due to the limitations in the resolution and the interpolation capabilities of current NeRF models, which may struggle to replicate the exact internal textures and structures needed for precise medical diagnosis and treatment planning.
(3) Ambiguity of Object Boundary. Medical images frequently run into ambiguity with object boundaries between different types of tissues or between pathological and healthy tissues. This ambiguity in boundary delineation poses a significant challenge for NeRFs, as these models rely on clear distinctions in visual data to reconstruct 3D spaces accurately. In medical imaging, where precise boundary identification is crucial for tasks such as tumor localization or the measurement of organ volumes, NeRFs must overcome this hurdle to ensure the reliability and accuracy of generated 3D models.
(4) Significance of Color Density. In medical imaging, variations in color density can convey critical information, such as the presence of blood flow (See Figure 2), variations in tissue density, or the accumulation of contrast agents. These subtle variations are essential for diagnosing conditions and planning treatments. NeRFs, which inherently model the light transport in a scene, must be adapted to interpret and accurately reproduce these color density variations in the synthesized 3D images. The challenge lies in the fact that NeRFs were initially designed for visual scenes where lighting plays a significant role in the appearance of objects, rather than for medical scenarios where color density variations are more closely related to organ properties and physiological conditions.
@@ 3 Methods
The unique properties of various organ structures introduce significant challenges to the reconstruction process, such as sparsity, overlap, and visibility issues, which can hinder the creation of complete morphological models, potentially resulting in missing details in the 3D reconstructions.
A prime example is the vessel structure, especially the challenge of vessel sparsity, where the thin nature of vessels requires a thorough sampling methodology for accurate depiction. Consequently, this section delves into a comprehensive classification based on the application of NeRFs across different organs. Our goal is to acquaint researchers with the essential operations and functionalities of NeRF models in medical imaging, addressing the intricacies involved in capturing the full complexity of organ structures.
@@ 3.1 NeRFs for Different Organs
Knee. Knee injuries and degenerative conditions, such as osteoarthritis, are very common [10]. Improved imaging and 3D reconstruction techniques can greatly aid in the diagnosis, treatment planning, and monitoring of such conditions, making the knee a highly relevant subject for research and application of NeRF technology. Furthermore, the knee is a common site for surgical interventions, including arthroscopy, ligament reconstruction, and total knee arthroplasty. NeRF-generated 3D models can enhance preoperative planning and simulation, providing surgeons with detailed visualizations of the anatomy, which can improve surgical outcomes.
@@
Recently, MedNeRF [5], a GAN-based Raditonal Filed, successfully untangles the surface shape, volumetric depth, and internal anatomical structures from 2D images by learning a function that assigns a radiance value to each pixel. MedNeRF is a Deep Learning model capable of reconstructing CT projections from limited or single-view X-ray images. Leveraging a novel architecture based on neural radiance fields, this model learns a continuous representation of CT scans. It effectively disentangles the shape and volumetric depth of both surface and internal anatomical structures from 2D images. However, due to the nature of GAN network architecture for image generation and using a single X-ray for CT projection reconstruction during image generation, the network may overfit to the single X-ray, leading to an imbalance between image clarity and accuracy. Thus, during the single-view volumetric rendering stage, MedNeRF fine-tuned the generator using the input view by jointly optimizing the reconstruction and the generator. The model showcases high-fidelity renderings, both qualitatively and quantitatively, and is compared favorably against other recent methods based on radiance fields, highlighting its effectiveness in reducing radiation exposure while maintaining imaging quality.
However, the hyperparameters in MedNeRF for balancing two losses (clarity and accuracy) are typically configured manually and subject to observation, and heavily depend on human expertise, which can pose challenges in achieving the optimal trade-off between image blurriness and accuracy. Manual adjustments of these loss weights make it hard to achieve the right balance between image blurriness and accuracy. To address this challenge, Uncertainty-aware MedNeRF (UMedNeRF) [6] is presented to automatically learn hyperparameters for clarity and precision balancing problems. As shown in Figure 3, the key to improving the quality of the generated images lies in simultaneously ensuring image resolution and accuracy. In that case, it becomes a multi-task learning problem. UMedNeRF uses an uncertainty-aware conditional radiance field that simultaneously weighs multiple loss functions automatically in the multi-task settings and generates the predicted patch for single-view volumetric rendering.
Fig. 4: Overview of extended-MedNeRF with 2D Slices of Brain MRI scans in [11].
More recently, alignment and correction NeRF (ACNeRF) [12] has been proposed to improve MedNeRF in two key areas: (1) enhancing the match between NeRF models and medical imagery by leveraging the models’ generalization capabilities more effectively, and (2) refining the reasoning process to better accommodate variations in rendering ranges. ACNeRF is designed to reconstruct projections of various poses from the projection of a single pose, it employs latent codes to linearly influence positional encoding, facilitating image alignment while ensuring the 3D structure remains intact by preventing generator parameter changes from causing disruptions. During the inference phase, a pose judgment network aids in correcting poses, thereby optimizing the model’s rendering views. For rendering narrow-range projections, ACNeRF incorporates a novel refinement approach that uses frequency-domain masks to fine-tune the generator.
Brain. Brain Magnetic Resonance Imaging (MRI) is pivotal in medical diagnostics and research because it allows for the detailed visualization of brain structures, aiding in the diagnosis, monitoring, and treatment planning of neurological conditions. Utilizing advanced imaging techniques, e.g., NeRFs, can enhance the quality and accuracy of brain images, and thus improve the detection of abnormalities, such as tumors, lesions, and degenerative diseases, providing crucial information for clinical decision-making and patient care. Besides, NeRFs can contribute to the development of more precise and less invasive diagnostic methods, improving patient outcomes and advancing the understanding of brain anatomy and function.
Specifically, the advent of NeRFs offers a promising solution, enabling the generation of precise 3D reconstructions from MRI images with minimal user input. Leveraging 3D convolutional neural networks, a recent work [11] reconstructs 3D projections from 2D MRI slices, overcoming challenges such as variable slice thickness and integrating multiple MRI modalities for enhanced robustness (see Figure 4). This technique facilitates accurate slice interpolation, capturing the shape and volumetric depth of both surface and internal structures. It also improves the comprehensiveness and reliability of 3D anatomical reconstructions, offering medical professionals the ability to visualize and analyze complex structures with fewer scans, potentially decreasing imaging times and mitigating motion-related complications.
Fig. 5: Overview of Masked NeRF [13]. Left: Skull projections from NeRF. Right. Skull projections from Masked NeRF scene representations.
Skulls. Skull Cone Beam Computed Tomography (CBCT) scans are the gold standard for diagnosing cranial fractures. They provide high-resolution images of the bone, allowing clinicians to identify even minor fractures that could be critical to a patient’s treatment plan. In the realm of image-guided, minimally invasive medical procedures, accurate pose estimation plays a pivotal role, utilizing X-ray projections to precisely navigate to targets within 3D space.
An unconstrained NeRF model applied to skull CBCT reconstructions exhibited artifacts beyond the reconstructed anatomical boundaries [13]. These artifacts compensate for overfitting to a constrained set of ground-truth X-ray images, resulting in implausible visualizations for novel poses. To mitigate these issues, a Masked-NeRF (mNeRF) approach is introduced by leveraging a 3D head mask derived from CBCT scans to spatially constrain NeRF’s extrapolations. The mask, deliberately expanded by 3mm for comprehensive coverage, ensures that densities beyond its perimeter are negated, facilitating seamless integration within the deep learning framework for gradient-based optimization and processing (See Figure 5).
Fig. 6: Overview of Coronary Angiography [14].
Heart Vessel. Imaging of heart vessels is crucial for diagnosing various heart conditions, including coronary artery disease (CAD), aneurysms, and blockages [15]. It allows doctors to assess the health of the heart’s blood vessels and identify areas of reduced blood flow. An example of Coronary Angiography is shown in Figure 6.
Conversely, the inherent characteristics of vessel structures present additional challenges, such as sparsity, overlap, and visibility, which obstruct the comprehensive reconstruction of vascular morphologies, potentially leading to omissions in the 3D models. These challenges include (1) vessel sparsity, which refers to the slender nature of vessels necessitating a comprehensive sampling approach for accurate reconstruction, (2) vessel overlap, concerning the concealment of crucial vessel segments in 2D projections due to occlusions, and (3) vessel visibility issues, where parts of the vessel may be obscured due to contrast variability or the superimposition of background structures.
A recent study [16] is dedicated to addressing the challenges inherent in the 3D reconstruction of coronary angiography images and exploring the applicability of the Neural Radiance Fields (NeRF) in overcoming these obstacles. Existing methods for the 3D reconstruction of coronary arteries often fall short, hampered by issues such as vessel morphology characteristics including sparsity, overlap, and distinguishing between foreground and background, alongside constraints posed by sparse view and limited angle X-ray projections. NeRF, with its promising outcomes in medical imaging (X-ray) applications, could address these challenges, although its effectiveness in handling specific nuances of X-ray angiography and vessel morphology has yet to be fully demonstrated. This work delves into the potential and limitations of employing NeRF for 3D reconstruction from X-ray angiography, conducting thorough experimental analyses to assess how angiographic challenges impact reconstruction quality. The results indicate NeRF’s promising capabilities for 3D reconstruction in the context of sparse and limited angle X-ray projections, while also pinpointing specific limitations, such as background overlap, that necessitate further research. Note that the method for large vessel Aorta is introduced in [17].
@@
Fig. 7: Overview of Sparse-view Neural Attenuation Fields (SNAF) [18]. An X-ray source systematically traverses around the region of interest, producing sparse projections captured by a flat-panel detector positioned on the opposite side.
Teeth. Teeth Cone Beam Computed Tomography (CBCT) images enable the detection and diagnosis of various dental conditions such as cavities, tooth decay, impacted teeth, and gum disease. They provide detailed views that are not visible to the naked eye, allowing for early treatment.
Neural Attenuation Fields (NAF) [17] innovatively models the attenuation coefficients as a continuous function across 3D space, utilizing a dense neural network for parameterization. NAF synthesizes projections and refines the network by reducing the discrepancy between actual and generated projections. A novel learning-based encoder, incorporating hash coding, enhances the model’s ability to delineate high-frequency details more effectively than traditional frequency-domain encoders. This improvement is attributed to the encoder’s ability to leverage the inherent smoothness and sparsity of human anatomical structures, leading to superior performance and efficiency in capturing intricate organ details.
@@
Furthermore, efforts to reconstruct high-quality CBCT images from sparse-view 2D projections have been ongoing, yet current leading methods often yield images marred by artifacts and missing fine details (See Figure 7). Sparse-view Neural Attenuation Fields (SNAF) [18] is proposed for sparse-view CBCT reconstruction. The view augmentation strategy addresses the data insufficiency challenges posed by sparse input views.
@@
Chest. Chest CT imaging stands as a cornerstone of medical diagnostics due to its ability to non-invasively reveal critical information about major organs, its role in a wide array of clinical applications, and its continual evolution alongside imaging technology advancements. The chest houses critical organs such as the lungs, heart, and major blood vessels. Imaging these structures can help diagnose a wide range of conditions, from respiratory infections and lung cancer to heart disease and vascular disorders. Chest images provide valuable insights into the anatomy and potential abnormalities within the thoracic cavity. The works in the previous section [5, 6, 12, 17] also are evaluated using the chest CT images.

### v4

# Neural Radiance Fields in Medical Image: A Survey

@@ Abstract
Neural Radiance Fields (NeRF), as a pioneering technique in computer vision, offer great potential to revolutionize medical imaging by synthesizing three-dimensional representations from projected two-dimensional image data. However, they face unique challenges when applied to medical imaging. This paper presents a comprehensive examination of NeRF applications in medical imaging, highlighting four imminent challenges, including fundamental imaging principles, internal structural requirement, object boundary definition, and color density significance. We discuss existing methods applied to different organs and their associated limitations. We also review several datasets and evaluation metrics and propose promising directions for future research.
@@ 1 Introduction
Neural Radiance Fields (NeRFs) significantly enhance traditional 3D reconstruction methods in several key aspects [1]. Unlike conventional techniques that often rely on geometric or voxel-based approaches, NeRFs utilize deep learning to create highly detailed and photorealistic 3D models from a set of 2D images [2]. Compared with traditional reconstruction methods, NeRFs can render complex scenes with intricate light and shadow interplay, which traditional methods struggle to capture accurately [3]. They also excel in handling occlusions and view-dependent effects, producing more continuous and smoother representations of 3D spaces. Additionally, NeRFs require fewer manual interventions and assumptions about scene geometry, leading to more automated and scalable 3D reconstructions [4]. This capability is particularly advantageous in medical imaging, where precise and detailed visualization of anatomical structures is crucial, offering enhanced diagnostic and surgical planning tools over traditional methods [5].
@@
By employing deep learning techniques to synthesize highly detailed and photorealistic 3D models from sets of 2D medical images, such as CT scans or MRIs, NeRFs provide an unprecedented level of detail and accuracy in visualizing complex anatomical structures [1, 8]. This technique surpasses traditional 3D reconstruction methods by effectively handling intricate light interactions and view-dependent phenomena, thus providing clearer, more continuous representations of internal body structures [9]. Furthermore, NeRFs can potentially minimize the need for multiple or high-resolution scans that expose patients to excessive ionizing radiation [6, 10]. Although the NeRFs-based method has yielded surprising results in traditional 3D image reconstruction, it cannot be directly applied to medical image reconstruction due to the unique internal details of medical images [11, 4].
This paper aims to highlight the significant motivation and challenges associated with implementing NeRFs in medical imaging. We have systematically categorized existing methods based on organ types, acknowledging that the distinct characteristics of various organ structures, such as sparsity, overlap, and visibility, pose considerable hurdles to achieving comprehensive morphological reconstructions. These obstacles often lead to incomplete 3D models, potentially omitting crucial details (see Figure 1 for an example). Then, we explore the future research opportunities that could help overcome these obstacles and advancing the field.
The key contributions of this work are as follows:
  • We outline the key motivations and challenges associated with applying NeRFs to medical imaging, identifying the main hurdles that must be overcome.
@@
  • Furthermore, we explore potential future research avenues, emphasizing how NeRFs in medical imaging can revolutionize diagnostic processes, surgical planning, and medical education. This represents a significant advancement in medical visualization and patient care, promising to enhance the accuracy and efficacy of treatments.
The remaining content of this paper is organized as follows. After introducing the motivation and challenges of NeRFs in medical imaging in Section 2, we describe the medical image NeRF applications for different organs and discuss the limitations of the current methods in Section 3. In Sections 4, we summarize the public datasets for NeRFs in medical imaging. We discuss the future directions in Section 5 and conclude the paper in Section 6.
@@ 2.1 Motivation
(1) Minimizing Radiation Exposure. NeRFs in medical imaging can potentially reduce the necessity for multiple or high-resolution scans that subject patients to high levels of ionizing radiation. By generating 3D representations from a sparse collection of 2D images, these methods can decrease the frequency of exposures, lowering radiation doses and reducing the risks associated with radiation exposure. This technique not only reduces the risk of radiation-related health conditions but it also alleviates issues related to prolonged patient immobility during imaging sessions.
(2) Reducing Time Requirements. Compiling multiple scans into a coherent 3D representation is time-intensive, posing significant challenges in time-sensitive medical situations. NeRFs in medical imaging streamline this process by enabling the quick creation of 3D models from a limited number of images, thereby accelerating data collection and image reconstruction. This rapid processing is particularly advantageous in emergency medicine settings, where efficiency can directly impact patient care and outcomes.
(3) Lowering Imaging Costs. Achieving detailed 3D images with conventional CT scans requires multiple X-ray projections, each with a fine slice thickness, which can be costly. NeRFs in medical imaging address this issue by efficiently producing detailed 3D models from smaller images, potentially reducing the financial burden. The capability of NeRFs in medical imaging to deliver precise volumetric reconstructions with fewer data inputs suggests a decrease in the reliance on expensive imaging resources, offering a cost-effective alternative to traditional imaging methods.
(4) Reducing Motion-related Artifacts. Medical imaging reconstructions typically involve multiple scan acquisitions, causing lengthy sessions that necessitate patient stillness, thereby increasing the risk of motion artifacts. Specifically, Magnetic Resonance Imaging (MRI) sequences are highly susceptible to motion-related artifacts caused by breathing and cardiac pulsation, potentially leading to inaccuracies in 3D reconstruction. Such inaccuracies are problematic in clinical settings, where precise anatomical measurements are essential for accurate treatment planning and patient monitoring.
@@ 2.2 Challenges
(1) Different Imaging Principles. The fundamental imaging principles of consumer electronics and medical imaging are different. For example, while color-based (RGB) NeRF methods have seen considerable development, their direct application to X-ray imaging may yield suboptimal outcomes due to inherent differences between visible light and X-ray modalities. Unlike visible light imaging, which captures external features by reflecting off an object’s surface, X-ray imaging penetrates and attenuates the object, creating images that primarily reveal internal structures. These internal details are crucial for 3D X-ray reconstruction, highlighting the need for tailored NeRF algorithms to effectively accommodate the unique characteristics of X-ray imaging [12].
(2) Need for Detailed Inner Structures. Medical images often require high levels of detail to represent intricate inner structures of the body accurately. To be clinically useful, NeRFs, which synthesize 3D models from 2D images, must capture these details. However, accurately rendering the complex internal anatomy, such as the fine distinctions between different types of tissues or the subtle anomalies indicative of early-stage diseases, can be challenging. This is due to the limitations in the resolution and the interpolation capabilities of current NeRF models, which may struggle to replicate the exact internal textures and structures needed for precise medical diagnosis and treatment planning.
(3) Ambiguity of Object Boundary. Medical images frequently run into ambiguity in object boundaries between different types of tissues or between pathological and healthy tissues. This ambiguity in boundary delineation poses a significant challenge for NeRFs, as these models rely on clear distinctions in visual data to reconstruct 3D spaces accurately. In medical imaging, where precise boundary identification is crucial for tasks such as tumor localization or the measurement of organ volumes, NeRFs must overcome this hurdle to ensure the reliability and accuracy of generated 3D models.
(4) Significance of Color Density. In medical imaging, variations in color density can convey critical information, such as the presence of blood flow (see Figure 2), variations in tissue density, or the accumulation of contrast agents. These variations are essential for diagnosing conditions and planning treatments. NeRFs, which inherently model the light transport in a scene, must be adapted to interpret and accurately reproduce these color density variations in the synthesized 3D images. The challenge lies in the fact that NeRFs were initially designed for visual scenes where lighting plays a significant role in the appearance of objects, rather than for medical scenarios where color density variations are more closely related to organ properties and physiological conditions.
@@ 3 Methods
The unique properties of various organ structures introduce significant challenges to the reconstruction process, such as sparsity, overlap, and visibility issues [13], which can hinder the creation of complete morphological models, potentially resulting in missing details in the 3D reconstructions [14]. A prime example is the vessel structure, especially the challenge of vessel sparsity, where the thin nature of vessels requires a thorough sampling methodology for accurate depiction. Consequently, this section provides a comprehensive classification based on the application of NeRFs across different organs. Our goal is to acquaint researchers with the essential operations and functionalities of NeRF models in medical imaging, addressing the intricacies involved in capturing the full complexity of organ structures.
@@ 3.1 NeRFs for Different Organs
Knee. Knee injuries and degenerative conditions, such as osteoarthritis, are very common [15]. Improved imaging and 3D reconstruction techniques can greatly aid in the diagnosis, treatment planning, and monitoring of such conditions, making the knee a highly relevant subject for research and application of NeRF technology. Recently, MedNeRF [6], a GAN-based Radiational Filed, successfully untangles the surface shape, volumetric depth, and internal anatomical structures from 2D images by learning a function that assigns a radiance value to each pixel. However, due to the nature of GAN-based network architecture and the reliance on a single X-ray for CT projection reconstruction, the network may overfit to the single X-ray, resulting in an imbalance between image clarity and accuracy. Thus, during the single-view volumetric rendering stage, MedNeRF fine-tuned the generator using the input view by jointly optimizing the reconstruction and the generator. The model showcases high-fidelity renderings, both qualitatively and quantitatively, and is compared favorably against other recent methods based on radiance fields, highlighting its effectiveness in reducing radiation exposure while maintaining imaging quality [16]. However, the hyperparameters in MedNeRF for balancing two losses (clarity and accuracy) are typically configured manually and subject to observation, and heavily depend on human expertise, which can pose challenges in achieving the optimal trade-off between image blurriness and accuracy. To address this challenge, Uncertainty-aware MedNeRF (UMedNeRF) [7] is presented to automatically learn hyperparameters for clarity and precision balancing problems. As shown in Figure 3, the key to improving the quality of the generated images lies in simultaneously ensuring image resolution and accuracy. In that case, it becomes a multi-task learning problem, where an uncertainty-aware conditional radiance field that simultaneously weighs multiple loss functions automatically in the multi-task settings and generates the predicted patch for single-view volumetric rendering.
@@
Fig. 4: (a) Overview of extended-MedNeRF with 2D Slices of Brain MRI scans in [17]. (b) Overview of Masked NeRF [18]. Left: Skull projections from NeRF. Right. Skull projections from Masked NeRF scene representations. (c) Overview of Coronary Angiography [19]. (d) Overview of Sparse-view Neural Attenuation Fields (SNAF) [20]. An X-ray source systematically traverses around the region of interest, producing sparse projections captured by a flat-panel detector positioned on the opposite side.
Brain. Brain Magnetic Resonance Imaging (MRI) is pivotal in medical diagnostics and research because it allows for the detailed visualization of brain structures, aiding in the diagnosis, monitoring, and treatment planning of neurological conditions. Utilizing advanced imaging techniques, e.g., NeRFs, can enhance the quality and accuracy of brain images, and thus improve the detection of abnormalities, such as tumors, lesions, and degenerative diseases, providing crucial information for clinical decision-making and patient care. NeRFs can contribute to the development of more precise and less invasive diagnostic methods, improving patient outcomes and advancing the understanding of brain anatomy and function. Specifically, the advent of NeRFs offers a promising solution, enabling the generation of precise 3D reconstructions from MRI images with minimal user input. Leveraging 3D convolutional neural networks, a recent study [17] reconstructs 3D projections from 2D MRI slices, overcoming challenges such as variable slice thickness and integrating multiple MRI modalities for enhanced robustness (see Figure 4 a). This technique facilitates accurate slice interpolation, capturing the shape and volumetric depth of both surface and internal structures. It also improves the accuracy and reliability of 3D anatomical reconstructions, offering medical professionals the ability to visualize and analyze complex structures with fewer scans, potentially decreasing imaging times and mitigating motion-related complications.
Skulls. Skull Cone Beam Computed Tomography (CBCT) scans are the gold standard for diagnosing cranial fractures. They provide high-resolution images of the bone, allowing clinicians to identify even minor fractures that could be critical to a patient’s treatment plan. In the realm of image-guided, minimally invasive medical procedures, accurate pose estimation plays a pivotal role, utilizing X-ray projections to precisely navigate to targets within 3D space. An unconstrained NeRF model applied to skull CBCT reconstructions exhibited artifacts beyond the reconstructed anatomical boundaries [18]. These artifacts stem from overfitting to a constrained set of ground-truth X-ray images, resulting in implausible visualizations for novel poses. To mitigate these issues, a Masked-NeRF (mNeRF) approach is introduced by leveraging a 3D head mask derived from CBCT scans to spatially constrain NeRF’s extrapolations. The mask, deliberately expanded by 3mm for comprehensive coverage, ensures that densities beyond its perimeter are negated, facilitating seamless integration within the deep learning framework for gradient-based optimization and processing (see Figure 4 b).
Heart Vessel. Cardiac vessel imaging is crucial for diagnosing various heart conditions, including coronary artery disease (CAD), aneurysms, and blockages [21, 22]. It allows doctors to assess the health of the heart’s blood vessels and identify areas of reduced blood flow. An example of coronary angiography is shown in Figure 4 c. However, the inherent characteristics of vessel structures present additional challenges, such as sparsity, overlap, and visibility, which obstruct the comprehensive reconstruction of vascular morphologies, potentially leading to omissions in the 3D models. These challenges include (1) vessel sparsity, which refers to the slender nature of vessels necessitating a comprehensive sampling approach for accurate reconstruction, (2) vessel overlap, concerning the concealment of crucial vessel segments in 2D projections due to occlusions, and (3) vessel visibility issues, where parts of the vessel may be obscured due to contrast variability or the superimposition of background structures. A recent study [23] is dedicated to addressing the challenges inherent in the 3D reconstruction of coronary angiography images and exploring the applicability of the NeRF in overcoming these obstacles. Existing methods for the 3D reconstruction of coronary arteries often fall short, hampered by issues such as vessel morphology characteristics including sparsity, overlap, and distinguishing between foreground and background, alongside constraints posed by sparse-view and limited-angle X-ray projections. The results indicate NeRF’s promising capabilities for 3D reconstruction in the context of sparse and limited angle X-ray projections, while also pinpointing specific limitations, such as background overlap, that necessitate further research. Note that the method for large vessel Aorta is introduced in [24].
@@
Teeth. Dental Cone Beam Computed Tomography (CBCT) images enable the detection and diagnosis of various dental conditions such as cavities, tooth decay, and gum disease [25, 26, 27]. They provide detailed views that are not visible to the naked eye, allowing for early treatment. Neural Attenuation Fields (NAF) [24] innovatively model the attenuation coefficients as a continuous function across 3D space, utilizing a dense neural network for parameterization. A novel learning-based encoder, incorporating hash coding, enhances the model’s ability to delineate high-frequency details more effectively than traditional frequency-domain encoders. Furthermore, efforts to reconstruct high-quality CBCT images from sparse-view 2D projections have been ongoing, yet current leading methods often yield images marred by artifacts and missing fine details (see Figure 4 d). Sparse-view Neural Attenuation Fields (SNAF) [20] has been proposed for sparse-view CBCT reconstruction. The view augmentation strategy addresses the data insufficiency challenges posed by sparse input views.
@@
Chest. Chest CT imaging is a cornerstone of medical diagnostics due to its ability to non-invasively reveal critical information about major organs, its role in a wide array of clinical applications, and its continual evolution alongside imaging technology advancements. The chest contains critical organs such as the lungs, heart, and major blood vessels. Imaging these structures can help diagnose a wide range of conditions, from respiratory infections and lung cancer to heart disease and vascular disorders. Chest images provide valuable insights into the anatomy and potential abnormalities within the thoracic cavity. The works in the previous section [6, 7, 16, 24] are also evaluated using the chest CT images.